# Edge-guided Low-light Image Enhancement with Inertial Bregman Alternating Linearized Minimization

### v1

@@ Abstract
Prior-based methods for low-light image enhancement often face challenges in extracting available prior information from dim images. To overcome this limitation, we introduce a simple yet effective Retinex model with the proposed edge extraction prior. More specifically, we design an edge extraction network to capture the fine edge features from the low-light image directly. Building upon the Retinex theory, we decompose the low-light image into its illumination and reflectance components and introduce an edge-guided Retinex model for enhancing low-light images. To solve the proposed model, we propose a novel inertial Bregman alternating linearized minimization algorithm. This algorithm addresses the optimization problem associated with the edge-guided Retinex model, enabling effective enhancement of low-light images. Through rigorous theoretical analysis, we establish the convergence properties of the algorithm. Besides, we prove that the proposed algorithm converges to a stationary point of the problem through nonconvex optimization theory. Furthermore, extensive experiments are conducted on multiple real-world low-light image datasets to demonstrate the efficiency and superiority of the proposed scheme.
@@ I Introduction
Due to lighting and environmental conditions, images captured by the camera may be low-light, which negatively impacts post-processing tasks like segmentation, classification, and recognition. Therefore, it is crucial to improve the quality of low-light images. Note that the Retinex theory describes the low-light image as the reflectance and the illustration part [1, 2]. By enhancing the reflectance and adjusting the illustration at the same time, the norm-light image can be obtained [3, 4]. Traditionally, the prior information is used to help enhance the reflectance [5, 6]. However, the detailed information, such as edge information, contained in the low-light image is difficult to obtain directly. As shown in Fig. 1, observed image (a) with low contrast, one can hardly recognize there is a bicycle, which is clear in the enhanced image (e). Besides, the edge information is hard to capture entirely by the traditional method as shown in (b). Hence, resulting the unsatisfied results (d). In this paper, we propose to learn clear edge information directly from the observed dim image to enhance the quality of the low-light image.
(a) Input
(b) Edge
(c) Our edge
(d) Retinex with (b)
(e) Ours
Figure 1: Low-light image enhancement. (a) is the observed low-light image; (b) is the edge obtained from (a) by gradient and divergence operator; (c) is the edge obtained from (a) by the proposed scheme; (d) is the result of Retinex method with edge (b); (e) is the result of our method with our edge (c). We evaluate the enhanced results with the Google Vision API and the probability of our results has a Bicycle being judged significantly.
@@
Generally, a low-light image $S$ can be decomposed as the reflectance part $R$ and illumination part $L$ [7], i.e., $S=R\circ L$, where $R\in(0,1)$ and $L\in(0,\infty)$. Directly enhancing the corresponding $R$ and $L$ to reconstruct the desired image $\hat{S}$ may be challenged for both $R$ and $L$ are unknown. Hence a minimization model was proposed as
$\min_{R,L}\frac{\beta}{2}\|S-R\circ L\|^{2}+\frac{\alpha}{2}\Psi(L)+\Phi(R),$ (1)
where the first term is the data fidelity term, $\Phi(\cdot)$ and $\Psi(\cdot)$ are the prior regularizers for $R$ and $L$, respectively, $\beta$ and $\alpha$ are positive parameters to balance the different terms. Based on the Retinex model, many crafted prior were designed to tackle the low-light image enhancement task. To better preserve fine details, Gu et al. [8] utilized the modified total variational model to fractional order for low-light image enhancement. The ADMM with convergence guarantee was proposed by adding auxiliary variables to solve the model. However, when their model faces the non-uniformed light image, the enhanced image may be unsatisfied. To overcome this disadvantage, Ma et al. [9] improved the model with a weighted map without theoretical analysis. By extending the fractional-order gradient total variation, small-magnitude details are maintained in the reflectance and the illumination is more likely piecewise smooth.
However, lower contrast often indicates smaller gradient magnitudes. Accordingly, the edge information cannot be obtained directly from the observed image. Moreover, the computational complexity of these models is expensive for the product form of the unknown variables. To tackle this problem, the logarithmic domain can be considered. Let $s=\log S$, $l=\log L$, and $r^{\prime}=\log R$, then we have $s=l+r^{\prime}$. Note that $R\in(0,1)$, we let $r=-r^{\prime}>0$, then $l=s+r$. Hence Eq. (1) can be reformulated as
$\min_{r,l}\frac{\beta}{2}\|l-s-r\|^{2}+\frac{\alpha}{2}\Psi(l)+\Phi(r).$ (2)
@@
In [12], the authors attempted to manipulate the gradient magnitudes of the reflectance to boost the contrast. By multiplying the gradient of the reflectance part with a constant, a more clear edge information can be collected. Similarly, a dual reflectance estimation method was proposed in [13] with convergence analysis to improve the performance of the Retinex model. In [14], a weighted map was designed to adjust the reflectance. Moreover, an alternating proximal gradient algorithm was proposed with convergence analysis to solve their Retinex model. Following this line, the weighted fractional-order total variation and dark channel prior were applied in [19] to amplify the edge information. Inspiring the hybrid deep learning prior [20, 21], Wu et al. [15] improved this model by adding a denoiser to remove the noise in the image. Instead of learning the data-fitting directly as RetinexNet [16], Liu et al. [17] further replaced $\Psi(L)$ with two implicit regularizers. A similar idea was presented in [18]. These works have competitive results in enhancing low-light images, however, their model and algorithm lack the theoretical guarantee. Besides, from these outstanding works, we know that edge information is essential to the low-light image enhancement task.
As mentioned before, extracting the edge information directly from the observed low-light image turns into unsatisfying results. Note that deep learning-based methods have revealed prominent performance in image processing. We wondering if the deep learning-based methods can help us to improve the results of the variational Retinex model. Typically, the edge $G$ of an image $S$ can be found by
$G=\mbox{div}(S_{x},S_{y}),$ (3)
where $S_{x}=\frac{\nabla_{x}S}{\sqrt{1+\left|\nabla S\right|^{2}}}$, $S_{y}=\frac{\nabla_{y}S}{\sqrt{1+\left|\nabla S\right|^{2}}}$, and $\mbox{div}(\cdot)$ denotes the divergence operation. However, when it comes to low-light images, as shown in Fig. 1 (b), the edge information is unsatisfied. Note that the authors in [22] applied the deep learning-based module to generate an edge prediction network to help the image super-resolution task. With such edge prediction network, the edge information of the low-resolution image can be extracted easily. A similar idea was used in [23] and [24] for image denoising and in [25] for MRI reconstruction.
@@
Inspired by these works, we aim to explore the deep edge prediction network to extract the edge feature from the low-light image and integrate the learned edge prior to the Retinex model. For simplicity, we will focus on the Retinex model in the logarithmic domain. To solve our low-light image enhancement model, we propose a novel inertial Bregman alternating linearized minimization algorithm. The details of the proposed algorithm will be introduced in Section II-C. For a clear perception, we summarize the essential comparison aspects in Table I. The contributions of the proposed low-light image enhancement scheme can be summarized as follows.
@@
  • We propose an edge-guided prior for low-light image enhancement, where the latent fine edge information can be exacted directly from the observed low-light image.
@@
  • With the help of the latent edge information, we integrate the proposed edge-guided prior to the Retinex model in the logarithmic domain for low-light image enhancement.
@@
The remainder of this paper is organized as follows. Section II reviews the basic concept of the Retinex theory, the edge prior, and some preliminaries for the proposed algorithm. The proposed edge extraction, image enhancement model, and algorithm are displayed in Section III. In Section IV, the convergence properties of the proposed algorithm are established. Experimental results on different datasets are in Section V. Finally, the conclusions follow in Section VI.
@@ II Related Works
In this section, we will give a short recall of Retinex theory on image processing, the edge prior, and some preliminaries for the proposed algorithm.
II-A Retinex image enhancement
The goal of Retinex theory [1] is to describe how the human visual system processes color changes and interprets the color of objects/scenes under changing illumination circumstances. In other words, Retinex theory explains that an object will show different colors under different lights, but people can know what color the object should be through common sense and visual senses. Therefore, the observed $S$ under different light can be described as
$S=R\circ L,$ (4)
where $R$ is the reflectance of the surface, and $L$ is the illumination on the surface.
With the product of reflectance and illuminance of an image, the Retinex-based image processing system is constructed. Hence, we only need to find $R$ and $L$ of a low-light image and enhance the intensity. However, it is difficult to find $R$ and $L$, directly. Consequently, model (1) is proposed with different regularizers to find better $R$ and $L$. Meanwhile, the logarithmic domain model (2) is proposed to reduce the computation complexity. In this paper, we focus on the model in the logarithmic domain for low-light image enhancement.
II-B Edge prior
If we find the edge of an image, the most detailed information of the image can be restored [26]. The variational method to extract the edge of an image $S$ is to constrain the gradient $\nabla S$. The discrete gradient operator $\nabla:\mathbb{R}^{m\times n}\rightarrow\left(\mathbb{R}^{m\times n},\ \mathbb{R}^{m\times n}\right)$ is defined by
$\vspace{-0.15in}(\nabla S)_{i,j}=\left(\left(\nabla_{x}S\right)_{i,j},\left(\nabla_{y}S\right)_{i,j}\right)$
with
$\left(\nabla_{x}S\right)_{i,j}=\left\{\begin{array}[]{lll}S_{i+1,j}-S_{i,j},&\text{ if }&i<m\\ S_{1,j}-S_{m,j},&\text{ if }&i=m\end{array}\right.$
and
$\left(\nabla_{y}S\right)_{i,j}=\left\{\begin{array}[]{lll}S_{i,j+1}-S_{i,j},&\text{ if }&j<n\\ S_{i,1}-S_{i,n},&\text{ if }&j=n\end{array}\right.$
for $i=1,\ldots,m$ and $j=1,\ldots,n$, where $S_{i,j}$ represents the value of pixel $(i,j)$ in the image.
Usually, the total variation model (ROF) [27] was proposed to preserve the edge of an image as
$\|\nabla S\|_{1}=\sum_{i,j}\left|(\nabla S)_{i,j}\right|=\sum_{i,j}\sqrt{\left(\nabla_{x}S\right)_{i,j}^{2}+\left(\nabla_{y}S\right)_{i,j}^{2}}.$
@@ II-B Edge prior
From the above literature, the edge of the image has its advantages for different tasks. In this paper, we learn the edge by using the convolutional neural network to help the image enhancement task.
Figure 2: The flowchart of the proposed scheme. We use 485 paired images from the LOL dataset as the training sets. The edge information can be directly obtained by using the proposed Edge-Net. The yellow, green, and blue blocks are the convolution layer, multi-scale residual block, and bottleneck layer, respectively. Then we use the learned edge prior to helping the image Retinex task. Through the proposed convergent algorithm, the low-light image can be well enhanced.
II-C Preliminaries
@@
    $\partial f(x):=\{u\in\mathbb{R}^{n}\;|\;\exists\ x^{k}\rightarrow x,~{}{\rm s.t.}~{}f(x^{k})\rightarrow f(x)$ (5)
@@
    $f$ is called a KL function if $f$ satisfies the KL property at each point of dom$(\partial f)$.
@@ III-A Edge extraction
$\mathcal{L}=\|G(S)-\widehat{S}_{egde}\|_{1},$ (7)
where $G(\cdot)$ represents the proposed edge network, $G(S)$ denotes the trained edge, and $\widehat{S}_{egde}$ is the edge label. Here the truth edge label is obtained from Eq. (3). As shown in Fig.2, we capture labels from the corresponding bright-light images. By minimizing the loss function, we train the edge network that can generate the edge directly from the low-light image. The detailed training setting is described in Section V.
III-B Retinex image enhancement model
As mentioned before, the intensity of $S$ mainly depends on the amount of source illumination falling on the objects in the scene and the amount of illumination reflected by the objects in the scene. Based on these two components, we can introduce two functions: the illumination function $L$ and the reflection function $R$. These two functions are assumed to be combined as a product to form image $S$
$S=L\circ R.$ (8)
Due to the reflectivity, $R\in(0,1)$. For the illumination effect, $L\in(0,\infty)$. Then, we know that $L>S>0$. To tackle the product form, we convert it into the logarithmic domain following [10], i.e.,
$s=\log(S),\quad l=\log(L),\quad r^{\prime}=\log(R).$ (9)
Then we obtain
$s=l+r^{\prime}.$ (10)
Note that $10<R<$ 1, hence $r^{\prime}<0$. We set $r=-r^{\prime}>0$, then the above model becomes
$l=s+r.$ (11)
Based on the former assumption, we propose the following energy functional for Retinex
$E(r,l)=\|\nabla l\|_{1}+\frac{\alpha}{2}\|\nabla r-\nabla g\|^{2}+\frac{\beta}{2}\|l-s-r\|^{2},$ (12)
where $g=G(S)$ is the learned edge prior, $\alpha$ and $\beta$ are the positive parameters.
III-C Inertial Bregman Alternating Minimization Algorithm
To solve the proposed model (12), we consider a more general class of structured optimization problems that minimize the following function:

### v2

@@ Abstract
Prior-based methods for low-light image enhancement often face challenges in extracting available prior information from dim images. To overcome this limitation, we introduce a simple yet effective Retinex model with the proposed edge extraction prior. More specifically, we design an edge extraction network to capture the fine edge features from the low-light image directly. Building upon the Retinex theory, we decompose the low-light image into its illumination and reflectance components and introduce an edge-guided Retinex model for enhancing the low-light image. To solve the proposed model, we propose a novel inertial Bregman alternating linearized minimization algorithm. This algorithm addresses the optimization problem associated with the edge-guided Retinex model, enabling effective enhancement of the low-light image. Through rigorous theoretical analysis, we establish the convergence properties of the algorithm. Besides, we prove that the proposed algorithm converges to a stationary point of the problem through nonconvex optimization theory. Furthermore, extensive experiments are conducted on multiple syntheses and real-world low-light image datasets to demonstrate the efficiency and superiority of the proposed scheme.
@@ I Introduction
Due to lighting and environmental conditions, images captured by the camera may be low-light, which negatively impacts post-processing tasks like segmentation, classification, and recognition. Therefore, it is crucial to improve the quality of the low-light image. Note that the Retinex theory describes the low-light image as the reflectance and the illumination part [1]. By enhancing the reflectance and adjusting the illumination at the same time, the norm-light image can be obtained [2]. Generally, the regularization term like the edge prior [3] can be utilized to enhance the low-light image. Furthermore, the deep edge extraction module [4] can also be applied for low-light image enhancement. However, extracting detailed edge information from low-light images is challenging. As shown in Fig. 1, one can hardly recognize the content of the observed low-contrast image (a). Besides, the edge structure of (e) in the dark area is not clear. To handle this problem, we propose to learn an edge extractor to obtain the detailed edge structure directly from the low-light image. In comparison to the edge produced by the SMG [4] in (g), the edge restored by our proposed edge extractor more closely resembles the ground-truth (GT) edge (h). With the assistance of the proposed edge extractor, we develop a novel Retinex model, which results in our enhancement outcome (c) demonstrating superior performance.
(a) (a) Input (10.74/50.87)
(b) (b) SMG (19.55/84.10)
(c) (c) Ours (24.18/92.45)
(d) (d) GT
(e) (e) edge of (a)
(f) (f) edge of SMG
(g) (g) edge of Ours
(h) (h) GT
Figure 1: Image enhancement results with index PSNR (dB) and SSIM (%). (a) is the input low-light image from dataset SYD [5]; (b) and (c) are enhancement results by SMG [4] and ours; (d) is the ground truth; (e) is the edge of (a) extracted by the Laplace operator; (f) and (g) are corresponding edges generated by SMG [4] and ours; (h) is the ground truth.
@@
Based on the Retinex theory, a low-light image $S$ can be decomposed as the reflectance $R$ and illumination part $L$, i.e., $S=R\circ L$, where $R\in(0,1)$, $L\in(0,\infty)$, and $\circ$ is the element-wise production operator. Directly obtaining $R$ and $L$ to reconstruct the desired image $\hat{S}$ is challenged for both $R$ and $L$ are unknown. Therefore, a minimization model has been proposed as follows:
$\min_{R,L}\frac{\beta}{2}\|S-R\circ L\|^{2}+\frac{\alpha}{2}\Phi(R)+\Psi(L),$ (1)
where the first term is the data fidelity term, $\Phi(\cdot)$ and $\Psi(\cdot)$ are the prior regularizers for $R$ and $L$, respectively, $\beta$ and $\alpha$ are positive parameters. Therefore, we only need to find $R$ and $L$ in the model and hence enhance the intensity of $S$.
For model (1), many researchers are devoted to estimating a fine reflectance map. For example, a dual reflectance estimation method was proposed in [6] to improve the performance of the Retinex model. Following this line, the weighted fractional-order total variation and dark channel prior were applied in [7] to enhance the reflectance part. Wu et al. [8] replaced the traditional prior with a convolutional neural network-based implicit regularizer to restore the reflectance map. Furthermore, Liu et al. [9] applied two implicit regularizers to adaptively fit both reflectance and illumination priors. A similar idea was presented in [10]. These works have competitive results in enhancing low-light images, however, most of them lack the theoretical guarantee.
To better preserve fine details, Gu et al. [11] utilized the modified total variational model to fractional order for low-light image enhancement. The alternating direction method of multipliers (ADMM) with convergence guarantee was proposed by adding auxiliary variables to solve the model. However, when their model faces the non-uniformed light image, the enhanced image may be unsatisfied. To overcome this disadvantage, Ma et al. [12] improved the model with a weighted map without theoretical analysis. By extending the fractional-order gradient total variation, small-magnitude details are maintained in the reflectance and the illumination is more likely piecewise smooth. Since the edge of an image refers to the curves changing sharply, and gradient can represent that “change” in the image [13]. Most edge detection methods rely on the computation of image gradients. However, lower contrast often indicates smaller gradient magnitudes. Accordingly, the edge information cannot be obtained directly from the observed image. The authors in [14] attempted to manipulate the gradient magnitudes of the reflectance to boost the contrast. By multiplying the gradient of the reflectance part with a constant, more clear edge information can be collected. This strategy has also been employed in [15, 8] for edge extraction in the context of low-light image enhancement. Drawing from these exemplary studies, it is evident that edge information is essential in low-light image enhancement.
@@
As previously mentioned, directly extracting edge information from the observed low-light image often yields unsatisfactory results. Deep learning-based methods, on the other hand, have shown exceptional performance in image processing. We are wondering whether these deep learning-based methods can assist us in enhancing the results of the variational Retinex model. Note that the authors in [20] applied the deep learning-based module to generate an edge prediction network with a Laplace loss function to help the image super-resolution task. With such an edge prediction network, the edge information of the low-resolution image can be extracted easily. A similar idea was explored in [21] and [22] for image denoising and in [23] for MRI reconstruction. In the realm of low-light image enhancement, the authors in [4] proposed an end-to-end neural network that incorporates an edge detection module to guide low-light image enhancement. Similarly, an edge enhancement module was introduced within an edge-enhanced multi-exposure fusion network [24] to enhance low-light images. In [25], a low-light image enhancement neural network was built with an edge module. From these edge-guided deep learning methods, it is apparent that the edge information is significant to image processing. However, the lack of interpretation is still a notable limitation.
On the other hand, the computational complexity of model (1) is expensive for the product form of two unknown variables. To tackle this problem, the logarithmic domain can be considered. Let $s=\log S$, $l=\log L$, and $r^{\prime}=\log R$, then we have $s=l+r^{\prime}$. Note that $R\in(0,1)$, we let $r=-r^{\prime}>0$, then $l=s+r$. Hence Eq. (1) can be reformulated as
$\min_{r,l}\frac{\beta}{2}\|l-s-r\|^{2}+\frac{\alpha}{2}\Phi(r)+\Psi(l).$ (2)
@@
Inspired by these works, we aim to explore the deep edge prediction network to directly extract the edge feature from the low-light image and integrate the learned edge extractor into the Retinex model. For simplicity, we will focus on the Retinex model in the logarithmic domain. To solve our low-light image enhancement model, we propose a novel inertial Bregman alternating linearized minimization algorithm. The details of the proposed algorithm will be introduced in Section II-B. For a clear perception, we summarize the essential comparison aspects in Table I. The contributions of the proposed low-light image enhancement scheme can be summarized as follows.
@@
  • We propose an edge extractor for low-light image enhancement, where the latent fine edge structure can be exacted directly from the observed low-light image.
@@
  • With the help of the latent edge structure, we integrate the proposed edge extractor into the Retinex model in the logarithmic domain for low-light image enhancement.
@@
The remainder of this paper is organized as follows. Section II reviews the basic concept of the edge regularizer and some preliminaries for the proposed algorithm. The proposed edge extraction, image enhancement model, and algorithm are displayed in Section III. In Section IV, the convergence properties of the proposed algorithm are established. Experimental results on different datasets are in Section V. Finally, the conclusions follow in Section VI.
@@ II Related Works
In this section, we will briefly recall the edge regularizer and some preliminaries for the proposed algorithm.
II-A Edge regularizer
The restoration of the most detailed information in an image relies on a careful extraction of its edge information [26]. However, for low-contrast images, conventional edge detection often fails to extract satisfactory edges (see Fig. 1(e)). In this paper, we aim to derive a deep edge exactor that reflects the local edge strength.
The traditional variational method to extract edge information involves modifying the image gradient $\nabla S$. Following the strategy in [20], we first compute the normalized gradients
$S_{x}=\frac{\nabla_{x}S}{\sqrt{1+\left|\nabla S\right|^{2}}},\quad S_{y}=\frac{\nabla_{y}S}{\sqrt{1+\left|\nabla S\right|^{2}}},$ (3)
where the $\nabla_{x}S$ and $\nabla_{y}S$ are horizontal and vertical gradients of the image $S$. Such normalization (3) suppresses low contrast variations and enhances prominent edges. The continuous edge indicator function $G$ is defined with standard divergence operation $\operatorname{div}(\cdot)$ as
$G=\operatorname{div}(S_{x},S_{y}).$ (4)
The detailed design of our edge extractor can be found in Section III-A. In discrete form, $G$ is approximated by the finite-difference expression
$G(i,j)=\frac{S_{x}(i+1,j)-S_{x}(i,j)}{\Delta x}+\frac{S_{y}(i,j+1)-S_{y}(i,j)}{\Delta y},$
where $\Delta x$ and $\Delta y$ are the grid spacings in the $x$- and $y$-directions, respectively.
While an idealized 45-degree edge (e.g., $S(x,y)=x+y$) in a continuous domain would theoretically yield $G=0$ due to uniform normalized gradients, this scenario assumes infinite contrast, perfect linearity, and noise-free conditions, which are rarely met in practice. Real-world edges exhibit finite contrast, discretization effects, and irregularities (e.g., noise or curvature), which disrupt gradient uniformity and ensure non-zero divergence.
The total variation regularizer, pioneered by Rudin, Osher, and Fatemi [27], known as the ROF regularizer, was proposed to preserve the edge of an image, which is defined by
$\|\nabla S\|_{1}=\sum_{i,j}\left|(\nabla S)_{i,j}\right|_{2}=\sum_{i,j}\sqrt{\left(\nabla_{x}S\right)_{i,j}^{2}+\left(\nabla_{y}S\right)_{i,j}^{2}}.$ (5)
This isotropic version of the ROF regularizer is known as the TV regularizer. The anisotropic version, as described in [28], is given by
$\hskip-7.22743pt\|\nabla S\|_{1}=\sum_{i,j}\left|(\nabla S)_{i,j}\right|_{1}=\sum_{i,j}{|\left(\nabla_{x}S\right)_{i,j}|+|\left(\nabla_{y}S\right)_{i,j}}|.$ (6)
@@ II-A Edge regularizer
Based on the aforementioned literature, edge information is beneficial for a variety of tasks. In this paper, we employ a convolutional neural network to learn the edge extractor, aiding in the image enhancement task.
Figure 2: The flowchart of the proposed scheme. We use 485 paired images from the LOL dataset as the training set. The edge information can be directly obtained by using the proposed Edge-Net. The yellow, green, and blue blocks are the convolution layer, multi-scale residual block, and bottleneck layer, respectively. Then we use the learned edge extractor to help with the image enhancement task. Through the proposed convergent framework, the low-light image can be well enhanced.
II-B Preliminaries
@@
    $\partial f(x)=\{u\in\mathbb{R}^{n}\;|\;\exists\ x^{k}\rightarrow x,~{\rm s.t.}~f(x^{k})\rightarrow f(x)$ (7)
@@
[…]
@@ III-A Edge extraction
$\mathcal{L}=\|G(S)-\widehat{S}_{edge}\|_{1},$ (9)
where $G(\cdot)$ represents the proposed edge exactor, $G(S)$ denotes the learning-based edge, and $\widehat{S}_{edge}$ is the edge label obtained from Eq. (4). Note that the learned image edge $G(S)$ has the same dimension as the input image $S$. As shown in Fig.2, we capture labels from the corresponding bright-light images. By minimizing the loss function, we train the edge network that can generate the edge directly from the low-light image. The detailed training setting is described in Section V.
III-B The proposed Retinex image enhancement scheme
Following the setting of Eq. (2) in the logarithmic domain, we propose to minimize the following energy functional based on Retinex theory for low-light image enhancement:
$E(r,l)=\frac{\beta}{2}\|l-s-r\|^{2}+\frac{\alpha}{2}\Phi(r,g)+\Psi(l),$ (10)
where $g=G(S)$ is the learned edge extractor, $\Phi(r,g)=\|\nabla r-\nabla g\|^{2}$, $\Psi(l)$ is a regularizer, $\alpha$ and $\beta$ are the positive parameters. To solve the proposed model (10), we consider a more general class of structured optimization problems that minimize the following function: